OpenStudy (anonymous):

Find the x- and y-intercepts of the graph of the equation. 2x − y − 9 = 0

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

what is y-intercepts?? and x-intercepts. what do they mean??

OpenStudy (anonymous):

X intercept is where the line crosses the x (horizontal) axis of the graph. Y intercept is where the line crosses the y (vertical) axis of the graph.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good. so Y intercept algebraically is when x=0. lets try it. >>>>> 2x − y − 9 = 0 before that lets have the eq in in y=mx+b y=2x-9 when x=0 y=(0)x -9 y=-9 >>>>>>Y intercept=-9

OpenStudy (anonymous):

Y intercept=-9 so we have (0,-9)

OpenStudy (anonymous):

X intercept algebraically is when y=0. 0=2x-9 9=2x 9/2=x >>>>>>X intercept=9/2

OpenStudy (anonymous):

hell yeah, thank you

OpenStudy (anonymous):

rate my respond if you like!

OpenStudy (anonymous):

thanks